OpenStudy (anonymous):

Is 100x^2-1 prime???

OpenStudy (jbo11):

are you trying to solve for x?

OpenStudy (anonymous):

Factor completly

hartnn (hartnn):

ax^2 +bx+c is prime only if b^2-4ac is not a perfect square so, find b^2-4ac for your question

OpenStudy (anonymous):

Just yes or no would be fine

hartnn (hartnn):

compare 100x^2-1 with ax^2+bx+c a=100. b=0, c=-1 then b^2-4ac = 0 - 4(100)(-1) =400 is 400 a perfect square ? just say yes or no, would be fine

OpenStudy (anonymous):

Yes 400 is a perfect square

OpenStudy (haseeb96):

b^2-4ac given a=100 , b=0 and c=-1 now put in it 0^2-4(100)(-1) 0+400=400 it is not prime no;

hartnn (hartnn):

correct! then your quadratic expression is NOT prime

OpenStudy (anonymous):

i'll cut to the chase without giving the answer: yes it can be factored.

hartnn (hartnn):

any more doubts on this?

OpenStudy (anonymous):

How can it be factored then?

OpenStudy (anonymous):

difference of squares

OpenStudy (anonymous):